# PyramidKV: Dynamic KV Cache Compression based on Pyramidal Information Funneling

### v2

@@ Abstract
In this study, we investigate whether attention-based information flow inside large language models (LLMs) is aggregated through noticeable patterns for long context processing. Our observations reveal that LLMs aggregate information through Pyramidal Information Funneling where attention is scattering widely in lower layers, progressively consolidating within specific contexts, and ultimately focusing on critical tokens (a.k.a massive activation or attention sink) in higher layers. Motivated by these insights, we developed PyramidKV, a novel and effective KV cache compression method. This approach dynamically adjusts the KV cache size across different layers, allocating more cache in lower layers and less in higher ones, diverging from traditional methods that maintain a uniform KV cache size. Our experimental evaluations, utilizing the LongBench benchmark, show that PyramidKV matches the performance of models with a full KV cache while retaining only 12% of the KV cache, thus significantly reducing memory usage. In scenarios emphasizing memory efficiency, where only 0.7% of the KV cache is maintained, PyramidKV surpasses other KV cache compression techniques, achieving up to a 20.5 absolute accuracy improvement on TREC.
@@ 1 Introduction
Large language models (LLMs) [1, 27, 28, 18] are integral to various natural language processing applications, including dialogue systems [5], document summarization [9], and code completion [25]. These models have recently been scaled up to handle long contexts [11, 7, 36, 4], with GPT-4 processing up to 128K tokens and Gemini-pro-1.5 handling 1M tokens. However, scaling LLMs to extremely long contexts naturally leads to a significant delay due to the quadratic computation of attention over long contexts. A common solution to mitigate such inference delays involves caching the key and value states (KV) of previous tokens [29], with the trade-off of requiring extensive GPU memory storage. For instance, maintaining a KV cache for 100K tokens in LLaMA-2 7B requires over 50GB of memory, while a 2K context requires less than 1GB of memory [31].
@@
Figure 1: Illustration of  PyramidKV compared with existing KV cache compression methods. (a) Full KV has all tokens stored in the KV cache in each layer; cache size increases as the input length increases. (b) StreamingLLM [32] only keeps few initial tokens with a fixed cache size in each layer. (c) SnapKV [22] and H2O [34] keep a fixed cache size across Transformer layers, and their selection is based on the attention score. (d)  PyramidKV maintains pyramid-like cache sizes, allocating more cache budget to lower layers and less to higher layers. This approach to KV cache selection better aligns with the increasing attention sparsity observed in multi-layer Transformers (§3).
@@
To tackle these memory constraints, recent studies have explored the optimization of KV caching, including approaches such as low-rank decomposition of the KV cache [8] or pruning non-essential KV cache [34, 22, 12]. Notably, it has been shown that maintaining merely 20% of the KV cache can preserve a substantial level of performance [34]. Moreover, extreme compression of the KV cache for tasks of longer contexts (e.g., retrieval augmented generation or RAG for short) can drastically improve efficiency and further reduce resource use. However, questions about the universal applicability of these strategies across all layers of an LLM remain open. (1) Are these KV cache strategies applicable to all layers? (2) Is it computationally efficient to use the same KV cache size across layers as previous studies have done? These considerations suggest a need for an in-depth, more nuanced understanding of KV cache optimization in LLMs.
@@
To examine these questions, we aim to systematically investigate the design principles of the KV cache compression across different layers, specifically tailored to the behaviors of the attention mechanism. We first investigate how information flow is aggregated via attention mechanisms across different layers in multi-document question answering (QA), a classic task involving long contexts. Our analysis identifies a notable transition of attention distribution from a broad coverage of global contexts to a narrow focus of local tokens over layers in LLMs. This pattern suggests an aggregated information flow where information is initially gathered broadly and subsequently narrowed down to key tokens, epitomizing the massive attention phenomenon. Our findings provide unique insights beyond the previously documented “massive activation” [26] that very few activations exhibit significantly larger values than others when calculating multi-head attention in LLMs and “attention sink” [32] that keeping the KV of initial tokens will largely recover the performance of window attention.
@@
Building on these insights on how information flows are aggregated through a pyramid pattern, we design a novel and effective KV cache pruning approach that mirrors the geometric shape, named PyramidKV. As shown in Figure 1, unlike the fixed-and-same length KV cache pruning common in prior works [34, 12, 22], PyramidKV allocates more KV cache to the lower layers where information is more dispersed and each KV holds less information while reducing the KV cache in higher layers where information becomes concentrated in fewer key tokens. To the best of our knowledge, PyramidKV is the first KV cache compression method with varied cache retention across layers, tailoring cache amounts to the informational needs of each layer and paving the way for future research.
@@
We conducted comprehensive experiments on LongBench [2] using 17 datasets across various tasks and domains with two backbone models (LLaMa-3-8B and Mistral-7B [18]). Results indicate that PyramidKV maintains performance using just 12.0% KV cache and excels in extreme conditions with only 0.7% KV cache retained. Moreover, PyramidKV outperforms baseline models (H2O [34], SnapKV [22], StreamingLLM [32]) across all tested cache sizes (64, 96, 128, 256, 512), with its advantages most pronounced at smaller cache sizes. Additionally, in the Needle In A Haystack experiment, PyramidKV notably sustains the long-context comprehension of LLMs better than competing methods, contributing to its superior performance in LongBench.
@@ 2 Related Work
Interpretation of LLMs
Prior research has shown that attention matrices in LLMs are typically sparse [3, 32, 34], focusing disproportionately on a few tokens. For instance, Xiao et al. [32] identified an “attention sink” phenomenon, where maintaining the Key and Value (KV) states of the first few tokens can substantially restore the performance of windowed attention, despite these tokens not being semantically crucial. Similarly, Sun et al. [26] identified a “massive activations” pattern, where a minority of activations show significantly larger values than others within LLMs. Interestingly, these values remain relatively constant across different inputs and act as critical bias terms in the model.
KV Cache Compression
There has been a growing interest in addressing LLMs’ memory constraints on processing long context inputs. FastGen [12] introduces an adaptive KV cache management strategy that optimizes memory use by tailoring retention tactics to the specific nature of attention heads. This method involves evicting long-range contexts from heads that prioritize local interactions, discarding non-special tokens from heads focused on special tokens, and maintaining a standard KV cache for heads that engage broadly across tokens. SnapKV [22] improves efficiency by compressing KV caches via selecting/clustering significant KV positions based on their attention scores. Heavy Hitter Oracle (H2O) [34] implements a dynamic eviction policy that effectively balances the retention of recent and historically significant tokens, optimizing memory usage while preserving essential information. StreamingLLM [32] enables LLMs trained on finite attention windows to handle infinite sequence lengths without fine-tuning, thus expanding the models’ applicability to broader contexts. LM-Infinite [15] allows LLMs pre-trained with 2K or 4K-long segments to generalize to up to 200M length inputs while retaining perplexity without parameter updates.
@@ 3 Pyramidal Information Funneling
In particular, we focus on our analysis of the LLaMa [27, 28] and visualize the distribution and behavior of attention scores over layers. To assess the distinct behaviors of each multi-head self-attention layer, we compute the average attention from all heads within each layer. Figure 2 shows the attention patterns of one QA example over six different layers (i.e., 0, 6, 12, 18, 24, and 30).
@@
Figure 2: Attention patterns of retrieval-augmented generation across layers in LlaMa [27, 28] reveal that in the lower layers, the model exhibits a broad-spectrum mode of attention, distributing attention scores uniformly across all content. In the middle layers, attention becomes more localized within each document, indicating refined information aggregation (dotted red triangular shapes in layers 6 and 10). This culminates in the upper layers, where “massive attention” focuses on a few key tokens (concentrated attention bars after layer 18), efficiently extracting essential information for answers.
@@ 4.1 Preliminaries and Problem Formulation
In an autoregressive transformer-based LLM, the generation of the $i$-th token requires that the attention module computes the query, key, and value vectors for all previous $i-1$ tokens. To speed up the inference process and avoid duplicate computations, the key and value matrices are typically stored in the GPU memory. While the KV cache enhances inference speed and reduces redundant computations, it can consume significant memory when dealing with long input contexts. To optimize memory usage, a strategy called KV cache compression is proposed [34, 32, 22], which involves retaining only a minimal amount of KV cache while preserving as much information as possible.
In a language model with $m$ transformer layers, we denote the key and value matrices in the $l$-th attention layer respectively as ${\bm{K}}^{l},{\bm{V}}^{l}\in\mathbb{R}^{n\times d},\forall l\in[0,m-1]$ when encoding a sequence of $n$ tokens. The goal of KV cache compression is to seek two sub-matrices ${\bm{K}}^{l}_{s},{\bm{V}}^{l}_{s}\in\mathbb{R}^{k^{l}\times d}$ from the full matrices ${\bm{K}}^{l}$ and ${\bm{V}}^{l}$, given a cache budget $k^{l}<n$ for each layer $l\in[0,m-1]$ while maximizing performance preservation. That is, a language model with KV cache compression only uses ${\bm{K}}^{l}_{s}$ and ${\bm{V}}^{l}_{s}$ in the GPU memory for inference on a dataset ${\mathcal{D}}$, and obtains a similar result to a full model according to an evaluation scoring metric, i.e., $\text{score}({\bm{K}}^{l},{\bm{V}}^{l},{\mathcal{D}})\approx\text{score}({\bm{K}}^{l}_{s},{\bm{V}}^{l}_{s},{\mathcal{D}})$.
@@ 4.2.1 KV Cache Size/Budget Allocation
Previous work on KV cache compression [22, 34, 32] often allocates a fixed KV cache size across LLM layers. However, as our analysis in §3 demonstrates, attention patterns are not identical across different layers. Particularly dense attention is observed in the lower layers, and sparse attention in higher layers. Therefore, using a fixed KV cache size across different layers may lead to suboptimal performance. These approaches may retain many unimportant tokens in the higher layers of sparser attentions while potentially overlooking many crucial tokens in the lower layers of denser attentions.
@@
Following the common practice in KV cache compression [22, 32], we first retain the KV cache for the last $\alpha$ tokens of the input across all layers, as these tokens have been shown to contain the most immediate task-related information, where $\alpha$ is a hyperparameter, controlling the number of last few tokens being included in the KV cache. For simplicity, we call these tokens “instruction tokens”, which is also referred to as “local window” in previous literature [34, 22, 32].
Subsequently, given the remaining total cache budget $k^{\text{total}}=\sum_{l\in[0,m-1]}k^{l}$ that can be used over all transformer layers (noted as $m$), we first determine the cache sizes for the top and bottom layers, and use an arithmetic sequence to compute the cache sizes for the intermediate layers to form the pyramidal shape. The key intuition is to follow the attention pattern in aggregated information flow, reflecting a monotonically decreasing pattern of important tokens for attention from lower layers to upper layers. We allocate $k^{0}=(2\cdot k^{\text{total}})/m$ for the first bottom layer, and $k^{m-1}=k^{\text{total}}/(\beta\cdot m)$ for the top layer, where $\beta$ is a hyperparameter to adjust the pyramid’s shape. Once the cache sizes of the bottom and top layers are determined, the cache sizes for all intermediate layers are set according to an arithmetic sequence, defined as
$k^{l}=k^{m-1}-\frac{k^{m-1}-k^{0}}{m}\times l.$ (1)
@@ 4.2.2 KV Cache Selection
Once the KV cache budget is determined for each layer, our method needs to select specific KV states for caching within each layer in LLMs. As described in the previous section, the KV cache of the last $\alpha$ tokens, referred to as instruction tokens, are retained across all layers. Following SnapKV [22], the selection of the remaining tokens is then guided by the attention scores derived from these instruction tokens—tokens receiving higher attention scores are deemed more relevant to the generation process and are thus their KV states are prioritized for retention in the GPU cache.
@@
where $d_{k}$ denotes the dimension of the key vectors. Following [22], we utilize a pooling layer at ${\bm{A}}^{h}$ to avoid the risk of being misled by some massive activation scores.
@@ 5 Experiment
We conduct comprehensive experiments to evaluate the effectiveness of PyramidKV on performance preserving and memory reduction. First, we introduce the backbone LLMs (§5.1), the evaluation datasets ( §5.2), and the baselines in comparison (§5.3). Next, we report the performance in a memory-oriented scenario and a performance-oriented scenario experiments in §5.5. We also test if the model can preserve the performance on long-context inputs on the Needle-in-the-haystack experiment ( §5.6). Finally, we discuss the trade-off between memory, time, and performance in §5.7.
5.1 Backbone LLMs
We compare PyramidKV against baselines using state-of-the-art open-sourced LLMs, namely LLaMa-3-8B-Instruct and Mistral-7B-Instruct [18]. Testing examples are evaluated in a generative format, with answers generated by greedy decoding across all tasks to ensure a fair comparison.
5.2 Datasets
We use LongBench[2] to assess the performance of PyramidKV on tasks involving long-context inputs. LongBench is a meticulously designed benchmark suite that tests the capabilities of language models in handling extended documents and complex information sequences. This benchmark was created for multi-task evaluation of long context inputs. It includes 17 datasets covering tasks such as single-document QA [20, 6], multi-document QA [33, 16], summarization [17, 35, 10], few-shot learning [21, 13, 19], synthetic, and code generation [14, 24]. The datasets feature an average input length ranging from 1,235 to 18,409 tokens (detailed average lengths can be found in Table 1), necessitating substantial memory for KV cache management. For all these tasks, we adhered to the standard metrics recommended by LongBench [2] (i.e., F1 for QA, Rouge-L for summarization, Acc. for synthetic and Edit Sim. for code generation.) We refer readers to more details at  Appendix B.
5.3 Baselines

### v4

@@ Abstract
In this study, we investigate whether attention-based information flow inside large language models (LLMs) is aggregated through noticeable patterns for long context processing. Our observations reveal that LLMs aggregate information through Pyramidal Information Funneling where attention is scattering widely in lower layers, progressively consolidating within specific contexts, and ultimately focusing on critical tokens (a.k.a massive activation or attention sink) in higher layers. Motivated by these insights, we developed PyramidKV, a novel and effective KV cache compression method. This approach dynamically adjusts the KV cache size across different layers, allocating more cache in lower layers and less in higher ones, diverging from traditional methods that maintain a uniform KV cache size. Our experimental evaluations, utilizing the LongBench benchmark, show that PyramidKV matches the performance of models with a full KV cache while retaining only 12% of the KV cache, thus significantly reducing memory usage. In scenarios emphasizing memory efficiency, where only 0.7% of the KV cache is maintained, PyramidKV surpasses other KV cache compression techniques, achieving up to a 20.5 absolute accuracy improvement on TREC dataset. In the Needle-in-a-Haystack experiment, PyramidKV outperforms competing methods in maintaining long-context comprehension in LLMs; notably, retaining just 128 KV cache entries enables the LLAMA-3-70B model to achieve 100.0 Acc. performance.
@@ 1 Introduction
Large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023a; b; Jiang et al., 2023) are integral to various natural language processing applications, including dialogue systems (Chiang et al., 2023), document summarization (Fabbri et al., 2019a), and code completion (Roziere et al., 2023). These models have recently been scaled up to handle long contexts (Fu et al., 2024; Ding et al., 2024; Zhu et al., 2023; Chen et al., 2023), with GPT-4 processing up to 128K tokens and Gemini-pro-1.5 handling 1M tokens. However, scaling LLMs to extremely long contexts naturally leads to a significant delay due to the quadratic computation of attention over long contexts. A common solution to mitigate such inference delays involves caching the key and value states (KV) of previous tokens (Waddington et al., 2013), with the trade-off of requiring extensive GPU memory storage. For instance, maintaining a KV cache for 100K tokens in LLaMA-2 7B requires over 50GB of memory, while a 2K context requires less than 1GB of memory (Wu et al., 2024).
@@
Figure 1: Illustration of PyramidKV compared with existing KV cache compression methods. (a) Full KV has all tokens stored in the KV cache in each layer; cache size increases as the input length increases. (b) StreamingLLM (Xiao et al., 2023) only keeps few initial tokens with a fixed cache size in each layer. (c) SnapKV (Li et al., 2024) and H2O (Zhang et al., 2024) keep a fixed cache size across Transformer layers, and their selection is based on the attention score. (d) PyramidKV maintains pyramid-like cache sizes, allocating more cache budget to lower layers and less to higher layers. This approach to KV cache selection better aligns with the increasing attention sparsity observed in multi-layer Transformers (§3).
@@
To tackle these memory constraints, recent studies have explored the optimization of KV caching, including approaches such as low-rank decomposition of the KV cache (Dong et al., 2024) or pruning non-essential KV cache (Zhang et al., 2024; Li et al., 2024; Ge et al., 2023). Notably, it has been shown that maintaining merely 20% of the KV cache can preserve a substantial level of performance (Zhang et al., 2024). Moreover, extreme compression of the KV cache for tasks of longer contexts (e.g., retrieval augmented generation or RAG for short) can drastically improve efficiency and further reduce resource use. However, questions about the universal applicability of these strategies across all layers of an LLM remain open. (1) Are these KV cache strategies applicable to all layers? (2) Is it computationally efficient to use the same KV cache size across layers as previous studies have done? These considerations suggest a need for an in-depth, more nuanced understanding of KV cache optimization in LLMs.
@@
To examine these questions, we aim to systematically investigate the design principles of the KV cache compression across different layers, specifically tailored to the behaviors of the attention mechanism. We first investigate how information flow is aggregated via attention mechanisms across different layers in multi-document question answering (QA), a classic task involving long contexts. Our analysis identifies a notable transition of attention distribution from a broad coverage of global contexts to a narrow focus of local tokens over layers in LLMs. This pattern suggests an aggregated information flow where information is initially gathered broadly and subsequently narrowed down to key tokens, epitomizing the massive attention phenomenon. Our findings provide unique insights beyond the previously documented “massive activation” (Sun et al., 2024) that very few activations exhibit significantly larger values than others when calculating multi-head attention in LLMs and “attention sink” (Xiao et al., 2023) that keeping the KV of initial tokens will largely recover the performance of window attention.
@@
Building on these insights on how information flows are aggregated through a pyramid pattern, we design a novel and effective KV cache pruning approach that mirrors the geometric shape, named PyramidKV. As shown in Figure 1, unlike the fixed-and-same length KV cache pruning common in prior works (Zhang et al., 2024; Ge et al., 2023; Li et al., 2024), PyramidKV allocates more KV cache to the lower layers where information is more dispersed and each KV holds less information while reducing the KV cache in higher layers where information becomes concentrated in fewer key tokens. To the best of our knowledge, PyramidKV is the first KV cache compression method with varied cache retention across layers, tailoring cache amounts to the informational needs of each layer.
Figure 2: Attention patterns of retrieval-augmented generation across layers in LlaMa (Touvron et al., 2023a; b) reveal that in the lower layers, the model exhibits a broad-spectrum mode of attention, distributing attention scores uniformly across all content. In the middle layers, attention becomes more localized within each document, indicating refined information aggregation (dotted red triangular shapes in layers 6 and 10). This culminates in the upper layers, where “massive attention” focuses on a few key tokens (concentrated attention bars after layer 18), efficiently extracting essential information for answers.
@@
We conducted comprehensive experiments on LongBench (Bai et al., 2023) using 17 datasets across various tasks and domains with three backbone models (LLaMa-3-8B-Instruct, LLaMa-3-70B-Instruct and Mistral-7B (Jiang et al., 2023)). The results show that PyramidKV preserves performance using just 12.0% of the KV cache (KV Cache size = 2048) on the LongBench benchmark and significantly outperforms other methods in extreme conditions, retaining only 0.7% of the KV cache. Moreover, PyramidKV outperforms baseline models (H2O (Zhang et al., 2024), SnapKV (Li et al., 2024), StreamingLLM (Xiao et al., 2023)) across all tested cache sizes (64, 96, 128, 256), with its advantages most pronounced at smaller cache sizes. In the Needle In A Haystack experiment, PyramidKV effectively maintains the long-context comprehension in LLMs, outperforming than competing methods. Remarkably, with PyramidKV, retaining only 128 KV cache entries allows the LLaMa-3-70B-Instruct model to achieve 100.0 Acc. performance, matching the performance of a full KV cache.
@@ 2 Related Work
There has been a growing interest in addressing LLMs’ memory constraints on processing long context inputs. FastGen (Ge et al., 2023) introduces an adaptive KV cache management strategy that optimizes memory use by tailoring retention tactics to the specific nature of attention heads. SnapKV (Li et al., 2024) improves efficiency by compressing KV caches via selecting/clustering significant KV positions based on their attention scores. Heavy Hitter Oracle (H2O) (Zhang et al., 2024) implements a dynamic eviction policy that effectively balances the retention of recent and historically significant tokens, optimizing memory usage while preserving essential information. StreamingLLM (Xiao et al., 2023) enables LLMs trained on finite attention windows to handle infinite sequence lengths without fine-tuning, thus expanding the models’ applicability to broader contexts.
@@ 3 Pyramidal Information Funneling
In particular, we focus on our analysis of the LLaMa (Touvron et al., 2023a; b) and visualize the distribution and behavior of attention scores over layers. To assess the distinct behaviors of each multi-head self-attention layer, we compute the average attention from all heads within each layer. Figure 2 shows the attention patterns of one QA example over six different layers (i.e., 0, 6, 12, 18, 24, and 30).
@@ 4.1 Preliminaries and Problem Formulation
In an autoregressive transformer LLM, the generation of the $i$-th token requires that the attention module computes the query, key, and value vectors for all previous $i-1$ tokens. To speed up inference process and avoid duplicate computations, the key and value matrices are typically stored in the GPU memory. While the KV cache enhances inference speed and reduces redundant computations, it can consume significant memory when dealing with long input contexts. To optimize memory usage, a strategy called KV cache compression is proposed (Zhang et al., 2024; Xiao et al., 2023; Li et al., 2024), which involves retaining only a minimal amount of KV cache while preserving as much information as possible.
In a LLM with $m$ transformer layers, we denote the key and value matrices in the $l$-th attention layer respectively as ${\bm{K}}^{l},{\bm{V}}^{l}\in\mathbb{R}^{n\times d},\forall l\in[0,m-1]$ when encoding a sequence of $n$ tokens. The goal of KV cache compression is to seek two sub-matrices ${\bm{K}}^{l}_{s},{\bm{V}}^{l}_{s}\in\mathbb{R}^{k^{l}\times d}$ from the full matrices ${\bm{K}}^{l}$ and ${\bm{V}}^{l}$, given a cache budget $k^{l}<n$ for each layer $l\in[0,m-1]$ while maximizing performance preservation. A LLM with KV cache compression only uses ${\bm{K}}^{l}_{s}$ and ${\bm{V}}^{l}_{s}$ in the GPU memory for inference on a dataset ${\mathcal{D}}$, and obtains a similar result to a full model according to an evaluation scoring metric, i.e., $\text{score}({\bm{K}}^{l},{\bm{V}}^{l},{\mathcal{D}})\approx\text{score}({\bm{K}}^{l}_{s},{\bm{V}}^{l}_{s},{\mathcal{D}})$.
@@ 4.2.1 KV Cache Size/Budget Allocation
Previous work on KV cache compression (Li et al., 2024; Zhang et al., 2024; Xiao et al., 2023) often allocates a fixed KV cache size across LLM layers. However, as our analysis in §3 demonstrates, attention patterns are not identical across different layers. Particularly dense attention is observed in the lower layers, and sparse attention in higher layers. Therefore, using a fixed KV cache size across layers may lead to suboptimal performance. These approaches may retain many unimportant tokens in the higher layers of sparser attentions while potentially overlooking many crucial tokens in the lower layers of denser attentions.
@@
Following the common practice in KV cache compression (Li et al., 2024; Xiao et al., 2023), we first retain the KV cache for the last $\alpha$ tokens of the input across all layers, as these tokens have been shown to contain the most immediate task-related information, where $\alpha$ is a hyperparameter, controlling the number of last few tokens being included in the KV cache. For simplicity, we call these tokens “instruction tokens”, which is also referred to as “local window” in previous literature (Zhang et al., 2024; Li et al., 2024; Xiao et al., 2023).
Subsequently, given the remaining total cache budget $k^{\text{total}}=\sum_{l\in[0,m-1]}k^{l}$ that can be used over all transformer layers (noted as $m$), we first determine the cache sizes for the top and bottom layers, and use an arithmetic sequence to compute the cache sizes for the intermediate layers to form the pyramidal shape. The key intuition is to follow the attention pattern in aggregated information flow, reflecting a monotonically decreasing pattern of important tokens for attention from lower layers to upper layers. We allocate $k^{m-1}=k^{\text{total}}/(\beta\cdot m)$ for the top layer and $k^{0}=(2\cdot k^{\text{total}})/m-k^{m-1}$ for the bottom layer,, where $\beta$ is a hyperparameter to adjust the pyramid’s shape. The hyperparameter $\beta$ is still required to determine the top layer. Once the top layer is identified, the budget of the bottom layer can be calculated by summing the budgets across all layers and equating this sum to the total budget. Once the cache sizes of the bottom and top layers are determined, the cache sizes for all intermediate layers are set according to an arithmetic sequence, defined as
$k^{l}=k^{0}-\frac{k^{0}-k^{m-1}}{m-1}\times l.$ (1)
@@ 4.2.2 KV Cache Selection
Once the KV cache budget is determined for each layer, our method needs to select specific KV states for caching within each layer in LLMs. As described in the previous section, the KV cache of the last $\alpha$ tokens, referred to as instruction tokens, are retained across all layers. Following SnapKV (Li et al., 2024), the selection of the remaining tokens is then guided by the attention scores derived from these instruction tokens—tokens receiving higher attention scores are deemed more relevant to the generation process and are thus their KV states are prioritized for retention in the GPU cache.
@@
where $d_{k}$ denotes the dimension of the key vectors. Following (Li et al., 2024), we utilize a pooling layer at ${\bm{A}}^{h}$ to avoid the risk of being misled by some massive activation scores.
@@ 5 Experiment
We conduct comprehensive experiments to evaluate the effectiveness of PyramidKV on performance preserving and memory reduction.
5.1 Experiment Setup
We maintain a fixed constant KV cache size for each layer for the baseline methods. In contrast, PyramidKV employs varying KV cache sizes across different layers. To ensure a fair comparison, we adjusted the average KV cache size in PyramidKV to match that of the baseline models, to keep the total memory consumption of all methods the same. We set $\beta=20$ and $\alpha=8$. We use the same prompt for each dataset in all experiments.
5.1.1 Backbone LLMs
We compare PyramidKV against baselines using state-of-the-art open-sourced LLMs, namely LLaMa-3-8B-Instruct, Mistral-7B-Instruct (Jiang et al., 2023) and LLaMa-3-70B-Instruct. Testing examples are evaluated in a generative format, with answers generated by greedy decoding across all tasks to ensure a fair comparison.
5.1.2 Datasets
We use LongBench (Bai et al., 2023) to assess the performance of PyramidKV on tasks involving long-context inputs. LongBench is a meticulously designed benchmark suite that tests the capabilities of language models in handling extended documents and complex information sequences. This benchmark was created for comprehensive multi-task evaluation of long context inputs. It includes 17 datasets covering tasks such as single-document QA (Kočiskỳ et al., 2018; Dasigi et al., 2021), multi-document QA (Yang et al., 2018; Ho et al., 2020), summarization (Huang et al., 2021; Zhong et al., 2021; Fabbri et al., 2019b), few-shot learning (Li and Roth, 2002; Gliwa et al., 2019; Joshi et al., 2017), synthetic, and code generation (Guo et al., 2023; Liu et al., 2023b). The datasets feature an average input length ranging from 1,235 to 18,409 tokens (detailed average lengths can be found in Table 1), necessitating substantial memory for KV cache management. For all these tasks, we adhered to the standard metrics recommended by LongBench (Bai et al., 2023) (i.e., F1 for QA, Rouge-L for summarization, Acc. for synthetic and Edit Sim. for code generation.) We refer readers to more details at Appendix F.
5.1.3 Baselines